Latest News







Site map




Rebuilding | Week 1 | Week 2/3 | Week 4 | Week 5 | Week 6 | Week 7 | Week 8 | Week 10 | Misc Thoughts

Monday, February 3, 2003

Over the weekend I picked up at 12v variable voltage regulated power supply and put together a new driver for the D&S system.

I spent Saturday putting the circuit together in between builds of the game I'm working on and was able to verify that the new driver at least drives the laser from 6v on up to 12v. It starts to dim visibly as the voltage goes below about 6v.

On Sunday I grabbed enough equipment to set up an interferometer, took it to work with me and did some voltage and temperature tests to see if the "fuzzing" out of the fringes changed at different temperatures and voltages.

What I found was that after the laser had gone through a 30min warm up and stabilized, I needed to lower the voltage to 7.5v to keep the fuzzies out of the fringe pattern. This also had the side effect of the laser putting out very high contrast fringes rather than the fuzzy-edged ones I'd become used to seeing from this laser.

During the day, I noticed that there were occasional very light fuzzies (my name for when the fringes blur into each other) and I began to wonder if air currents were causing them. I found some clear stiff plastic wrap and created a tube to shield the laser head from most currents.

That seemed to take care of the fuzzies I was seeing.

Rather than going to sleep once I got home (1:30am) I of course set everything up on my table and ran the tests again.

While the tests results were similar, I was consistently getting more fuzzies at home than at work.

I began to wonder if I was seeing additional problems with air currents at home vs. work because I could blow on the mounts and watch the fringes fuzz out in a way that was similar to what I'd see if I didn't blow at the table. This I found interesting because I never saw that behavior at work, where the mounts were merely sitting on my desk, which is non-magnetic.

Am I seeing contractions of the metal top (1/4" steel) due to temperature shifts or am I just seeing currents cross the table?

At this point I'm still not sure what the cause is, but since I'm seeing the behavior with two different drivers (same design) I believe I can rule out the driver, unless I'm putting the thing together wrong.

One of the many other things I'm going to add to the tests is a way to easily insert by HeNe into the interferometer layout so I can sanity check the two lasers against each other.


This evening I mounted my HeNe on a tripod, and used it in the interferometer setup.

Rock solid, high contrast fringes (once I'd left the room and it had settled for about 20s).

I then put the diode laser onto a second tripod and inserted it into the same interferometer setup.

After settling, rock solid, high contrast fringes with no fuzzing. Ever.

Now I'm really confused. Is the laser itself (the driver is actually still sitting on the table, with wires running to the diode head) causing the fuzzing out I'm seeing? Is there some sort of resonance between it, the mount it's on and the table?

I now know that I've got two other problems (especially after reading various archives from this and other boards) with the table top and the inner tubes.

The top is clearly not rigid enough. I can easily cause it to flex as I move my magnetic mounts around and pushing slowly on any part of the table causes the fringes to flow like water.

The inner tubes may be too full. I need to get the stem extensions put on so I can easily modify the pressure until the top floats better.

Wednesday, February 5, 2003

Rigging a remote shutter and switch so that I can trigger the exposure from across the room.  This is a temporary measure until I come up with a way to connect the shutter to either my PC or a palm pilot.  I'm thinking about writing some custom timing software for the pilot that can control the shutter and give off timing beeps during development.

CAT5 cable connected to a small normally-open switch

with the other end connected to my main power supply and the closed shutter.  The clear plastic is a temporary measure to reduce the impact of air currents near the laser head.

Another view of the shutter in use and open.

Thursday, February 6, 2003

Well, turns out I was partly wrong concerning the past failures.. The reflection holograms I've been making were (more or less) working. The problem is that they're playing back too far into the red and are almost invisible in white light.

This morning I changed three things:

1) Changed the subject to a small aluminum model of the Millennium Falcon. The first model was a "walking" Snoopy and it occurred to me that the internal spring may be causing movement of the model.

2) Knowing that they are less sensitive to vibration problems I changed the setup for a transmission hologram.

3) Moved the plate holder so that it was attached to the table rather than the right-angle mount (they both use magnets). This way, the plate holder was more secure.

The transmission worked the first time although imo it's a bit dim. There's no picture because I still haven't been able to get a decent one with the digital camera.

Next I changed the setup to make a reflection version of the same scene.

That one came out as a dud. Or, at least, I thought it did. When trying to find the image using a flashlight I noticed that there was *something* dim and red there. For no reason I can think of, I decided to take a look at it in laser light and there it was, the Millennium Falcon behind a piece of plastic (fairly bright too).

Now I've just got to figure out if I need to pre-swell the film before exposure or do some extra post processing or do something else entirely to get it to play back in green.

I then took the other holograms I'd made and looked at them in laser light. I could clearly see black rings on Snoopy's belly and nose, which indicate that he was in motion. I think the internal spring is exerting enough pressure on him that he's constantly moving.

The laser is still showing odd behavior in the interferometer and I'll continue to try to find the source of that problem but at least I now know that the general setup and processing is working.

Sunday, February 9, 2003

Before going in to work today I ran two tests using the Millennium Falcon setup.

1) After an exposure of 3s and development (but before bleaching) I dipped the hologram into a cup of fixer diluted 1:3 with water.  After 5 seconds I lifted the film out about 1/4" and repeated this process until I ran out of film.

2) I tripled the exposure and skipped the fixer step.

In the first test (hologram #11), there was no apparent change in the hologram.  It was only visible in laser light and none of the fixed areas was in any way color shifted.  I may try this test later with a more concentrated fixer solution.

In the second test, the general image was brighter but still only visible in laser light.  The image also showed regular bands in the plane of the hologram (except where the mirror was reflecting the light back onto the model, and then they changed direction) that seem to indicate the diode laser was changing frequency.


Rebuilding | Week 1 | Week 2/3 | Week 4 | Week 5 | Week 6 | Week 7 | Week 8 | Week 10 | Misc Thoughts


Use this form to send us your questions or comments about this web site or the products featured here.

Copyright (c) 1996-2015 Dragon's Eye Software, All Rights Reserved.
All trademarks are owned by the respective company, Dragon's Eye Software or Dragon's Eye Creations

Last modified: Monday, April 21, 2008